OpenStudy (danielbarriosr1):

How many gallons of heavy cream at 9% butterfat should be mixed with milk 2% butterfat to obtain 35 gallons of milk with 4% butterfat?

OpenStudy (nikato):



OpenStudy (nikato):

so 0.09x+0.02(35-x)=0.04(35) solve for x

OpenStudy (anonymous):

x= - 61/2

OpenStudy (danielbarriosr1):

x=10